OpenStudy (anonymous):

Find the equations of both lines that pass through the origin and are tangent to the parabola y =x² + 4

OpenStudy (anonymous):

first lets call the points \[P(a_{1},b_{1}),Q(a,b)\]

OpenStudy (anonymous):

their gradient is \[f \prime (x)=2x\]

OpenStudy (anonymous):

since they are passing through the origin they have the equation \[y=2(a)x\] \[y=2(a _{1})x\]

OpenStudy (anonymous):

all we have to do now is find the points intersection with the graph

OpenStudy (anonymous):

\[a^2+4=2(a)a\]

OpenStudy (dumbcow):

lines tangent to parabola have slope of 2x y=mx+b since they must go through origin, b=0 --> y = (2x)*x = 2x^2 find point where x^2 +4 = 2x^2 --> x = +-2 therefore 2 lines are y = +-4x

OpenStudy (anonymous):

\[2a^2-a^2=4\] \[a^2=4\] \[x=\pm 2\]

OpenStudy (anonymous):

very nice dumbcow

OpenStudy (anonymous):

y cordinate\[y=2(2)=4\] \[y=2(-2)=-4\]